# Tel.: +91-080-2293 2058

### v1

@@ 3.3.2 Beltrami Flow
Figure 8: Beltrami Flow: PINTO solutions (first row), corresponding analytical solution (second row) and relative error of PINTO solution (fourth row) for seen (first two columns) and unseen (last two columns) initial conditions (Re) at solution time step $t=0.5$. Flow streamlines are overlaid on a background of the velocity magnitude $|V|=\sqrt{u^{2}+v^{2}}$ for the solution. Only magnitude of relative error is shown.
@@ 3.3.3 Lid Driven Cavity Flow
Table 1 documents the mean relative error in the predictions of the PINTO on the $64\times 64$ grid for the seen and unseen lid velocities. For validation, numerical solutions for all lid velocities during training and testing are simulated from the Finite Volume code (Ueckermann and Lermusiaux, 2012). Figure  10 compares PINTO and PI-DeepONet solutions for unseen lid velocities between 1 and 3 (the values seen during training) and greater than 3 (outside the range of values seen during training). For the lid velocity of $1.2$, the PINTO and PI-DeepONet solutions are accurate and comparable to the numerical solution, whereas, for the lid velocity of $3.5$, the PINTO solutions are better compared to the PI-DeepONet solutions. For PI-DeepONet, in this challenging test, the relative error crosses 12% at several grid points, making the solution unusable. In general, the relative error for different unseen lid velocities from the PINTO solution is $2.78\%$, which is less than half of the relative error ($6.08\%$) of the PI-DeepONet solution.
@@ 4 Discussion
Computational complexity: Generally, transformer operators are computationally intensive due to the attention matrix. In addition, transformer-based neural operators Hao et al. (2023) require extensive simulation or observational data to learn to generalize to irregular meshes and multiple input functions. In contrast, PINTO achieves generalization using cross-attention units trained by minimizing the physics-loss. As the query point is a sequence of length $M=1$, the complexity of PINTO’s attention mechanism is $O\left(N\times n_{e}^{2}\right)$, as opposed to $O\left(MN\times n_{e}^{2}\right)$ of general attention, where $M,N$ represents query and key sequence lengths and $n_{e}$ is the embedding dimension.
@@
$\min_{\theta}\,\frac{1}{N_{c}}\sum_{j=1}^{N_{c}}|f_{c}^{j}-\mathcal{N}(\mathcal{G}(\Theta;X_{c}^{j});\alpha)|^{2}+\frac{1}{N_{ib}}\sum_{j=1}^{N_{ib}}|b_{ib}^{j}-\mathcal{B}(\mathcal{G}(\Theta;X_{ib}^{j}))|^{2}\,,$ (12)
@@
where, $\{(X_{ib}^{j},b_{ib}^{j})\}_{j=1}^{N_{ib}}$ denote the training data points on $\partial\Omega$, and $\{X_{c}^{j}\}_{j=1}^{N_{c}}$ are the training points (called collocation points) on the domain $\Omega$. $\mathcal{G}(X;\Theta^{*})$ is an approximate solution to Eq. 1 for a given $b$. For a different $b$, the optimization problem in Eq. 12 must be solved again to obtain another set of optimal parameters $\left(\Theta^{*}\right)$. In other words, PINNs need retraining if the initial and boundary conditions change. A neural operator framework is necessary to generalize to different initial and boundary conditions. However, the neural operators in the literature are trained using simulation data, hence we proposed our neural operator that efficiently generalizes to different initial/boundary conditions while training with physics-loss.
@@
Interpretation from classical numerics: Neural operators are akin to traditional numerical techniques, such as finite elements, substituting the linear span of local basis functions with neural operator layers Li et al. (2020a). The initial and boundary conditions ensure the necessary constraints for a unique PDE solution, dictating the function space for the solution, so the network must learn to incorporate these conditions. Numerical methods involve discretizing the PDE Eq. 1 into a system of linear equations $(Ax=b)$, providing solutions across the grid. Alternatively, the point solution corresponds to the weighted sum of the initial and boundary conditions ($x=A^{-1}b$). Interior points as queries, with initial and boundary points as keys and values in an attention mechanism, yield the weighted sum of boundary functions, mirroring numerical methods. The attention score combines two terms: (i) similarity and (ii) magnitude Song et al. (2021). The similarity term uses an RBF kernel with a length-scale hyperparameter to assess the query-key importance. Thus, our PINTO model can be viewed as a learning scheme inspired by traditional numerical methods, to us the attention mechanism for learning boundary-aware vectors.
@@
Comparison to other neural operators: It is worth noting that traditional physics-informed neural networks or neural operators require separate training for each set of initial/boundary conditions, whereas our PINTO model allows training once for multiple initial/boundary conditions and can handle unseen conditions during inference. The physics-informed DeepONet (PI-DeepONet) (Lu et al., 2021a; Goswami et al., 2023) used in the present paper for the comparison of PINTO was repurposed to find PDE solutions with different initial/boundary conditions. Unlike PI-DeepONet, PINTO allows different vector lengths to be used as input of the BPE and BVE units during training and testing rather than a fixed length. By means of the cross-attention mechanism, the query point’s representation vector effectively captures all the relevant information from the boundary points and their values, which is necessary for predicting the PDE solution at that specific query point for any unseen boundary condition. As our goal was to train with physics-loss, other neural operator models such as physics-informed neural operators, Fourier neural operators, or convolutional neural operators that need simulation data for training at some resolution are not considered for comparison.
@@
Unlike the DeepONet architecture (Lu et al., 2021a; Wang et al., 2021), one key novelty of PINTO is the treatment of the initial/boundary conditions as a sequence of vectors, where each element is a vector of initial/boundary coordinates and values. In addition, PINTO uses an attention mechanism to learn the encoded representation of the initial/boundary functional space $\mathcal{B}$. This mode of representation not only helps in generalization but also gives the practical utility of using different lengths of boundary vectors during training and inference phases, making our approach further different from the branch and trunk nets (DeepONet).
@@
Limitations: Since the training of the PINTO model uses collocation points and physics-loss, it may suffer spectral bias reported in other PINN models. There might also be training discrepancies due to imbalances in different terms in the loss functions. Adaptive training and dynamic weighting strategies (Jin et al., 2021; Daw et al., 2023; Hou et al., 2023; Maddu et al., 2022; Deguchi and Asai, 2023; Tang et al., 2023; Moseley et al., 2023) must be adopted to address these limitations.
@@ 5 Conclusion and Future Scope
Our newly developed PINTO architecture enhances the generalizability of neural models to solve IBVPs with multiple initial and boundary conditions. The key novelty of the architecture is the introduction of the new cross-attention unit that implements the attention-based integral kernel operator. This architecture is successful in transforming the representation vector of the PDE solution’s domain to a boundary/initial condition-aware encoding. Using five diverse test cases, we demonstrated that the relative error of the PINTO model is comparable for seen and unseen initial and boundary conditions. These results show that the encoding learned through our cross-attention unit is efficient in generalizing to new out-of-distribution (unseen during training) input functions of initial and boundary conditions. Comparative analysis with the PI-DeepONet neural operator model demonstrates the superior performance of PINTO to effectively generalize to unseen initial and boundary conditions. The relative error of PINTO solutions for all test cases in unseen testing configurations is low, only 20% to 33% of the relative error of PI-DeepONet’s solutions. We also show that PINTO is able to find solutions outside the temporal domain in which training is performed, the first result of this nature in the neural operator literature. Future directions of our work include the incorporation of turbulent dynamics into the PINTO framework. We envision that the PINTO model will find applications in optimizing the layout of wind farms, building digital twins, modeling the earth system, and any other domain governed by initial boundary value problems. Moreover, the cross-attention units are versatile, hence, they could be used to devise architectures for geometry generalization as well. Incorporating uncertainty quantification schemes Guo et al. (2024); Zou et al. (2024) into PINTO and using it for inverse problems are also potential future directions Gao et al. (2024); Guo et al. (2022).
Appendix A Choice of hyperparameters:
@@
Like any neural model, the performance of PINTO is influenced by the number of layers, units, the activation function, and other hyperparameters. Here, the decision of the number of cross-attention units (CAUs) in the second stage of PINTO, the sequence length of initial and boundary conditions, and the choice of training hyperparameters are made based on cross validation. Since here we are emphasizing the applicability of our proposed architecture to generalize for unseen initial and boundary conditions, we have chosen the hyperparameters that gave good predictions with low relative errors on the unseen conditions; alternatively, one can choose the optimal set of hyperparameters using the grid search algorithm or efficient search algorithms (Bergstra et al., 2011; Escapil-Inchauspé and Ruz, 2023; Kaplarević-Mališić et al., 2023; Wang and Zhong, 2024). Table 2 summarizes the hyperparameters selected for Advection and Burgers equations. For solving Navier Stokes equation, PINTO has learned the mapping with lesser relative errors (Table. 1) with a single CAU unit and shallow multi-layer perceptrons for lifting and projection units with activation function of Swish. Due to high imbalances in the boundary values for lid-driven cavity flow, while training both PINTO and PI-DeepONet, the loss term with 00 boundary value has given a weight of $100$.
@@
Appendix B Hyperparameters and Learning Curves
Learning curves during training of PINTO and PI-DeepONet for the five test cases are shown in Figure 11. For a fair comparison between the PINTO and PI-DeepONet solutions, the models are built approximately with the same number of trainable parameters and have used similar training strategies to ensure their convergence.

### v3

@@ 3.3.2 Beltrami Flow
Figure 9: Beltrami Flow: PINTO solutions (first row), corresponding analytical solution (second row) and relative error of PINTO solution (fourth row) for seen (first two columns) and unseen (last two columns) initial conditions (Re) at solution time step $t=0.5$. Flow streamlines are overlaid on a background of the velocity magnitude $|V|=\sqrt{u^{2}+v^{2}}$ for the solution. As the error is small, $\log_{10}$ of magnitude of the relative error is shown so that it is visible.
@@
Figure 10: Mean relative error of PINTO predictions for the Kovasznay and Beltrami flows at different Reynolds numbers.
Figure 10 shows the performance of PINTO at different Reynolds numbers (Re). For the Kovasznay flow, the mean relative error for a given Re is computed across all discrete points in the spatial domain. For the Beltrami flow, for a given Re, the mean relative error is computed across all discrete spatio-temporal points in the spatio-temporal domain. The minimum and maximum Re considered to train the PINTO model are 20, 80208020,\,8020 , 80 for the Kovasznay and 10, $0,\,10$ 100 for the Beltrami Flow. In both cases, as Re increases, there is an increase in the mean relative error, but it is less than 1.5% even at very large unseen Re that were not shown during the training. These results demonstrate that PINTO training is robust and generalizable.
@@ 3.3.3 Lid Driven Cavity Flow
Table 2 documents the mean relative error in the predictions of the PINTO on the $64\times 64$ grid for the seen and unseen lid velocities. For validation, numerical solutions for all lid velocities during training and testing are simulated from the Finite Volume code [83]. Figure  12 compares PINTO and PI-DeepONet solutions for unseen lid velocities between 1 and 3 (the values seen during training) and greater than 3 (outside the range of values seen during training). For the lid velocity of $1.2$, the PINTO and PI-DeepONet solutions are accurate and comparable to the numerical solution, whereas, for the lid velocity of $3.5$, the PINTO solutions are better compared to the PI-DeepONet solutions. For PI-DeepONet, in this challenging test, the relative error crosses 12% at several grid points, making the solution unusable. In general, the relative error for different unseen lid velocities from the PINTO solution is $2.78\%$, which is less than half of the relative error ($6.08\%$) of the PI-DeepONet solution.
@@ 4 Discussion
Computational complexity: Generally, transformer operators are computationally intensive due to the attention matrix. In addition, transformer-based neural operators [55] require extensive simulation or observational data to learn to generalize to irregular meshes and multiple input functions. In contrast, PINTO achieves generalization using cross-attention units trained by minimizing the physics-loss. As the query point is a sequence of length $M=1$, the complexity of PINTO’s attention mechanism is $O\left(N\times m^{2}\right)$, as opposed to $O\left(MN\times m^{2}\right)$ of general attention, where $M,N$ represents query and key sequence lengths and $m$ is the embedding dimension (dimension of encoded vector). C shows the computational time for all test cases.
@@
$\min_{\theta}\,\frac{1}{N_{c}}\sum_{j=1}^{N_{c}}|f^{j}-\mathcal{N}(\mathcal{G}(\Theta;X^{j});\alpha)|^{2}+\frac{1}{N_{ib}}\sum_{j=1}^{N_{ib}}|b^{j}-\mathcal{B}(\mathcal{G}(\Theta;X_{b}^{j}))|^{2}\,,$ (12)
@@
where, $\{(X_{b}^{j},b^{j})\}_{j=1}^{N_{ib}}$ denote the training data points on $\partial\Omega$, and $\{X^{j}\}_{j=1}^{N_{c}}$ are the training points (called collocation points) on the domain $\Omega$. $\mathcal{G}(X;\Theta^{*})$ is an approximate solution to Eq. 1 for a given $b$. For a different $b$, the optimization problem in Eq. 12 must be solved again to obtain another set of optimal parameters $\left(\Theta^{*}\right)$. In other words, PINNs need retraining if the initial and boundary conditions change. A neural operator framework is necessary to generalize to different initial and boundary conditions. However, the neural operators in the literature are majorly trained using simulation data and recently with some works reporting physics-informed training. In the present work, our focus is on developing a neural operator that efficiently generalizes to different initial/boundary conditions while training only with physics loss.
@@
Interpretation from classical numerics: Neural operators are akin to traditional numerical techniques, such as finite elements, substituting the linear span of local basis functions with neural operator layers [6]. The initial and boundary conditions ensure the necessary constraints for a unique PDE solution, dictating the function space for the solution, so the network must learn to incorporate these conditions. Numerical methods involve discretizing the PDE Eq. 1 into a system of linear equations $(Ax=b)$, providing solutions across the grid. Alternatively, the point solution corresponds to the weighted sum of the initial and boundary conditions ($x=A^{-1}b$). Interior points as queries, with initial and boundary points as keys and values in an attention mechanism, yield the weighted sum of boundary functions, mirroring numerical methods. The attention score combines two terms: (i) similarity and (ii) magnitude [84]. The similarity term uses an RBF kernel with a length-scale hyperparameter to assess the query-key importance. Thus, our PINTO model can be viewed as a learning scheme inspired by traditional numerical methods, to use the attention mechanism for learning boundary-aware vectors. For the advection equation we show an explicit calculation in  D to illustrate the above interpretation.
@@
Comparison to other neural operators: It is worth noting that traditional physics-informed neural networks require separate training for each set of initial/boundary conditions, whereas our PINTO model allows training once for multiple initial/boundary conditions and can handle unseen conditions during inference. The physics-informed DeepONet (PI-DeepONet) [7, 36] used in the present paper for the comparison of PINTO was repurposed to find PDE solutions with different initial/boundary conditions. In DeepONets, the vector of input functions is passed through the branch network. The branch network is an MLP, with the input layer dimension equal to the length of the vector; since the shape of the weight matrix is fixed between the input layer and the first hidden dense layer, it is not possible to pass vectors of different lengths to the DeepONets during the inference phase. On the other hand, the attention mechanism processes the input as a sequence of tokens. Since PINTO uses an attention mechanism, it allows different vector lengths as input for the BPE and BVE units during training and testing rather than a fixed length without affecting the model architecture. By means of the cross-attention mechanism, the query point’s representation vector effectively captures all the relevant information from the boundary points and their values, which is necessary for predicting the PDE solution at that specific query point for any unseen boundary condition. As our goal was to train with physics-loss, other neural operator models such as physics-informed neural operators, Fourier neural operators, or convolutional neural operators that need simulation data for training at some resolution are not considered for comparison.
@@
Unlike the DeepONet architecture [7, 30], one key novelty of PINTO is the treatment of the initial/boundary conditions as a sequence of vectors, where each element is a vector of initial/boundary coordinates and values. In addition, PINTO uses an attention mechanism to learn the encoded representation of the initial/boundary functional space $\mathcal{B}$. This mode of representation not only helps in generalization but also gives the practical utility of using different lengths of boundary vectors during training and inference phases, making our approach further different from the branch and trunk nets (DeepONet).
@@
Extending generalization to other parameters: In this paper we emphasized generalization for any initial and boundary conditions. However, the idea of the cross attention unit can be extended to other types of generalization such as geometry or other flow parameters. For example, if the position of obstacles or shape of objects in a flow field can be parametrized to a functional representation, then PINTO could be applied to the parametrized sequence of geometry. For generalizing for unseen fluid properties (such as viscosity), the cross attention units can attend to the sequence of parametrized properties.
Extension to 3D in space test cases: The extension of PINTO to 3D in space is straightforward. The number of collocation points and boundary points will increase, leading to an increase in the training time and memory requirement. An increase in training instabilities due to vanishing or exploding gradients, and imbalances in the loss terms, could also arise. Furthermore, the complexity of the cross-attention units is more than PI-DeepONet, thus training will be more expensive in comparison.
Physics guided training and discrete loss functions: In this paper, we focus on physics-informed training using automatic differentiation. However, the proposed architecture can also be trained using simulation data in a physics-guided setting.  E shows results for the physics-guided training of Burger’s equation. Instead of automatic differentiation, discrete loss functions based on finite difference, finite volume [85, 86] and finite element [87, 16]
Limitations: Since the training of the PINTO model uses collocation points and physics-loss, it may suffer spectral bias, a common phenomenon while training neural networks, and has also been reported in other PINN models. There might also be training discrepancies due to imbalances in different terms in the loss functions. Adaptive training, adaptive sampling, and dynamic weighting strategies [88, 89, 90, 91, 92, 93, 94] must be adopted to address these limitations.
@@ 5 Conclusion and Future Scope
Our newly developed PINTO architecture enhances the generalizability of neural models to solve PDEs with multiple initial and boundary conditions. The key novelty of the architecture is the introduction of the new cross-attention unit that implements the attention-based integral kernel operator. This architecture is successful in transforming the representation vector of the PDE solution’s domain to a boundary/initial condition-aware encoding. Using five diverse test cases, we demonstrated that the relative error of the PINTO model is comparable for seen and unseen initial and boundary conditions. These results show that the encoding learned through our cross-attention unit is efficient in generalizing to new out-of-distribution (unseen during training) input functions of initial and boundary conditions. Comparative analysis with the PI-DeepONet neural operator model demonstrates the superior performance of PINTO to effectively generalize to unseen initial and boundary conditions. The relative error of PINTO solutions for all test cases in unseen testing configurations is low, only 20% to 33% of the relative error of PI-DeepONet’s solutions. We also show that PINTO is able to find solutions outside the temporal domain in which training is performed, the first result of this nature in the neural operator literature. Future directions of our work include the incorporation of turbulent dynamics into the PINTO framework. We envision that the PINTO model will find applications in optimizing the layout of wind farms, building digital twins, modeling the earth system, and any other domain governed by initial boundary value problems. Moreover, the cross-attention units are versatile, hence, they could be used to devise architectures for geometry generalization as well.
Acknowledgements
We thank the members of our QUEST Lab at the Indian Institute of Science Bangalore for useful discussions, especially Rishi Jinka and Azhar Gafoor. This work was partially funded by research grants from the Ministry of Earth Sciences (MoES/36/OOIS/Extra/84/2022).
Appendix A Performance metrics
In this appendix, we define the relative error metric shown in Section 3 and other error metrics, and report results to analyze the PINTO and PI-DeepONet models. Section 3 presents results exclusively for the adjusted relative error. This adjustment (Eq. 17) is implemented to handle scenarios in which true values are nearly zero; in these instances, the absolute error is employed. This approach enhances the interpretability of the model’s performance. The definitions of all error metrics used in the paper are as follows.
[…]
where, $N_{q}$ is total number of query points, $\hat{h}_{i}$ is the prediction of $i^{th}$ query point from the model and $h_{i}$ is the true value of $i^{th}$ query point.
Tables 3 and  4 highlight that the generalization for the PINTO model is better than the PI-DeepONets across multiple performance metrics.
Table 3: Different performance metrics for all test cases
Test Cases PINTO PI-DeepONets
Training Testing Training Testing
RMSE MAE RMSE MAE RMSE MAE RMSE MAE
Advection $0.042$ $0.026$ $0.074$ $0.0398$ $0.0243$ $0.0147$ $0.2185$ $0.1587$
Burgers $0.0469$ $0.0245$ $0.0421$ $0.0258$ $0.0692$ $0.0334$ $0.0834$ $0.0546$
Kovasznay $0.0009$ $0.0007$ $0.0186$ $0.0046$ $0.0022$ $0.0016$ $0.069$ $0.0338$
Beltrami $0.0138$ $0.0068$ $0.0157$ $0.0091$ $0.0598$ $0.0338$ $0.1616$ $0.0759$
Lid Driven $0.0307$ $0.0193$ $0.0642$ $0.045$ $0.0514$ $0.0288$ $0.1983$ $0.1055$
Table 4: Different performance metrics for all test cases
Test Cases PINTO PI-DeepONets
Training Testing Training Testing
nRMSE MAPE nRMSE MAPE nRMSE MAPE nRMSE MAPE
Advection $0.0585$ $0.0471$ $0.1037$ $0.0733$ $0.0334$ $0.02611$ $0.3049$ $0.2921$
Burgers $0.2738$ $0.1895$ $0.2711$ $0.2112$ $0.4044$ $0.2578$ $0.5376$ $0.4467$
Kovasznay $0.0007$ $0.0006$ $0.0114$ $0.0035$ $0.0017$ $0.00151$ $0.0573$ $0.0398$
Beltrami $0.0295$ $0.0171$ $0.0336$ $0.0227$ $0.1276$ $0.0848$ $0.2709$ $0.1529$
Lid Driven $0.0527$ $0.0494$ $0.077$ $0.073$ $0.0905$ $0.0739$ $0.2189$ $0.1708$
Appendix B Choice of hyperparameters
@@
Like any neural model, the performance of PINTO is influenced by the number of layers, units, the activation function, and other hyperparameters. Here, the decision of the number of cross-attention units (CAUs) in the second stage of PINTO, the sequence length of initial and boundary conditions, and the choice of training hyperparameters are made based on cross validation. Since here we are emphasizing the applicability of our proposed architecture to generalize for unseen initial and boundary conditions, we have chosen the hyperparameters that gave good predictions with low relative errors on the unseen conditions; alternatively, one can choose the optimal set of hyperparameters using the grid search algorithm or efficient search algorithms [95, 96, 97, 98]. Table 5 summarizes the hyperparameters selected for Advection and Burgers equations. For solving Navier Stokes equation, PINTO has learned the mapping with lesser relative errors (Table. 2) with a single CAU unit and shallow multi-layer perceptrons for lifting and projection units with activation function of Swish. Due to high imbalances in the boundary values for lid-driven cavity flow, while training both PINTO and PI-DeepONet, the loss term with 00 boundary value has given a weight of $100$.
@@
Hyperparameters and Learning Curves Learning curves during training of PINTO and PI-DeepONet for the five test cases are shown in Figure 13. For a fair comparison between the PINTO and PI-DeepONet solutions, the models are built approximately with the same number of trainable parameters and have used similar training strategies to ensure their convergence.
@@
Appendix C Computational Time
The Table 8 provides the inference time for each test case. The inference time is reported as the mean of $10$ individual forward passes performed on Nvidia 48GB RTX A6000 single GPU machine.
Table 8: Training and inference time
Test Case Inference Time Grid Size
in milli seconds (ms)
1D Advection equation $40$ms $1024\times 100\,\left(x,t\right)$
1D Burgers equation $35$ms $1024\times 100\,\left(x,t\right)$
Kovasznay Flow $185$ms $256\times 256\,\left(x,y\right)$
Beltrami Flow $275$ms $64\times 64\times 20\,\left(x,y,t\right)$
Lid Driven Cavity Flow $100$ms $256\times 256\,\left(x,y\right)$
Appendix D Interpretation from FDM
Consider the implicit discretization scheme with a forward difference in time and a center difference in space for the 1D advection equation (Eq. 9) with Dirichlet boundary conditions, and the expression to compute the solution at $\left(n+1\right)^{th}$ step from $n^{th}$ time step for spatial domain with $4-$ grid points is as follows,
$\begin{bmatrix}u_{0}^{n+1}\\ u_{1}^{n+1}\\ u_{2}^{n+1}\\ u_{3}^{n+1}\end{bmatrix}=\begin{bmatrix}1&0&0&0\\ \frac{\psi}{1+\psi^{2}}&\frac{1}{1+\psi^{2}}&0&\frac{\psi^{2}}{1+\psi^{2}}\\ \frac{\psi^{2}}{1+\psi^{2}}&\frac{\psi}{1+\psi^{2}}&\frac{1}{1+\psi^{2}}&\frac{-\psi}{1+\psi^{2}}\\ 0&0&0&1\\ \end{bmatrix}\begin{bmatrix}u_{0}^{n}\\ u_{1}^{n}\\ u_{2}^{n}\\ u_{3}^{n}\\ \end{bmatrix}.$
Here, $\psi=\frac{\beta\Delta t}{2\Delta x}$, $\Delta t$, and $\Delta x$ are discretization lengths in time and space respectively. The solution in the $\left(n+1\right)^{th}$ step of a grid point is a weighted sum of all grid points at $n^{th}$ time step; iteratively, it is a weighted sum of initial values and the weights are a function of $\Delta x$ and $\Delta t$. Now, consider the attention score $\zeta_{i}$ from eq. 7, neglecting the normalizing constant and replacing $A_{h}\mu_{t}\left(X\right)$ with $q$, and $B_{h}MLP_{bpe}\left(X_{b}^{i}\right)$ with $k_{i}$ the resulting expression will be $\exp\left(\frac{q^{T}k_{i}}{\sqrt{m}}\right)$, further it is decomposed into two terms [84] as follows:
$\exp\left(\frac{-||q-k_{i}||^{2}_{2}}{2\sqrt{m}}\right)\,\times\exp\left(\frac{||q||^{2}_{2}+||k_{i}||^{2}_{2}}{2\sqrt{m}}\right),$
the first term is the RBF kernel with fixed length scale hyperparameter, $4]{m}$ that measures the similarity score between query and key using squared Euclidean distance $||q-k_{i}||^{2}_{2}$, and the second term is the magnitude. The computation of attention scores has a nonlinear relationship with the distance between query and key, similar to the weights in the discretization scheme discussed above, which are functions of discretization lengths.
Appendix E Physics-Guided training for Burger’s equation test case
In this Appendix, we present results from simulation data-driven training (physics-guided neural transformer operator - PGNTO) of our architecture for the Burger’s equation test case. Here, we don’t use physics loss and compare results with the DeepONet. Table 9 shows that the PINTO model performs better than the DeepONet. Figure 14 shows the learning curves of PINTO and DeepONet during training.
Table 9: Hyperparameters and results of physics-guided training for the Burger’s equation test case
Model Epochs Optimizer Learning Trainable Training Performace Metrics
Rate Parameters Time Training Testing
PGNTO $5000$ Adam $41e-$ 4 $141825$ $75$mins $0.691\%\,\left(0.646\%\right)$ $1.26\%\,\left(1.52\%\right)$
DeepONets $5000$ Adam $41e-$ 4 $140288$ $5$mins $1.81\%\,\left(1.874\%\right)$ $3.01\%\,\left(2.82\%\right)$
Figure 14: Learning curves of PINTO and PI-DeepONets model for data-driven training of Burgers equation.